The Student Information Processing Board (SIPB)

• What is SIPB?

The Student Information Processing Board (SIPB, pronounced sip-bee) is the MIT volunteer student group dedicated to improving computing at MIT. We are a basically a group of MIT students, staff, and alumni who like working with computers and making useful contributions to the community.

• What does SIPB do to “improve computing”?

We interpret improving computing to mean a lot of different things. SIPB members answer a lot of walk-in and phone questions about computers. Some SIPB members maintain servers used by the MIT community, including the original MIT web server (www.mit.edu), Usenet news service (news.mit.edu), and an AFS cell (/afs/sipb.mit.edu). SIPB has produced an installer for Linux-Athena for personal workstations. Some members are working on a port of Athena to Debian Linux. We document inessential parts of Athena, such as account customization and zephyr. Many members work on outside coding projects including the Debian project, GNOME, OMS (a free player for DVD and other media) in addition to SIPB projects. SIPB members also teach IAP classes to MIT students on a variety of topics including networking, Perl, software development and system administration. You probably received this flyer in one such class. Most members just hang out in the office, hacking or tooling and answering walk-in or phone questions.

• Can I join SIPB?

Yes. Anyone in the MIT community can become a SIPB prospective. Just come by the office and ask. To become a member, first become a prospective. Then hang around, get to know people, and do useful stuff. The general idea is to show that you are interested in what SIPB does, that you will benefit SIPB. Once people know you, you will be put up for membership. Full members are able to allocate SIPB’s money, have keys to the office, and access to some other restricted resources.

• Why should I join SIPB?

The best reason to join SIPB is that you are interested in learning about and helping people with computers. Many SIPB members are very skilled; it is amazing what one can learn just by hanging around in the office. Members are generally willing to answer questions, talk about their projects, and help you with yours. SIPB has other resources, like hardware and money, which you may be able to use to help with a project. And if you are just bored and looking for something to hone your skills on, you can find a project that needs help.

Another good reason to become a SIPB member is to have a place to come and hack, talk and hang out with other computer geeks. The SIPB office is one place where people will never tell you that your conversation is too geeky.

• Where do I sign up?

The SIPB office is W20-557, just outside of the Student Center cluster. The office is open whenever a member is there, which is most of the 168 hours in a week. Feel free to drop by, talk to us, and ask about becoming a prospective. Official meetings are Monday at 7:30 in the office. They are good for knowing what’s going on in the SIPB. Right after a meeting there are usually more members and prospectives around than at other times, but meetings are still no substitute for spending time around the office in general.